The wave function of a quantum particle of mass m is πœ“ (π‘₯) = 𝐴 cos(π‘˜π‘₯) + 𝐡 sin(π‘˜π‘₯) where 𝐴, 𝐡 π‘Žπ‘›π‘‘ π‘˜ are constants. (a) Assuming the particle is free (π‘ˆ = 0), show that πœ“ (π‘₯) is a solution of the π‘†π‘β„Žπ‘ŸΓΆπ‘‘π‘–π‘›π‘”π‘’π‘Ÿ equation. (b) Find the corresponding energy E of the particle.


1
Expert's answer
2021-01-17T14:45:09-0500

The wave function of a quantum particle satisfies the Schrodinger equation

"-\frac{\hbar^2}{2m}\frac{d^2\psi(x)}{dx^2}+U(x)\psi(x)=E\psi(x)"

In the case of free particle

"-\frac{\hbar^2}{2m}\frac{d^2\psi(x)}{dx^2}=E\psi(x)""\frac{d^2\psi(x)}{dx^2}=\frac{d^2}{dx^2}\left(A\cos(kx)+B\sin(kx)\right)\\
=-k^2\left(A\cos(kx)+B\sin(kx)\right)=-k^2\psi(x)"

Hence, the "\psi(x)" is a solution of the Schrodinger equation if


"E=\frac{\hbar^2 k^2}{2m}"